Can someone show me how to solve by the substitution method. Thanks!
4X - y = -3
y = 4X + 3

Answer by tutorcecilia(2144)

4X - y = -3
y = 4X + 3
.
substitute one variable within the other equation.
if y=4x+3, plug it in the other equation
.
4x-y=-3
4x-(4x+3)=-3 [solve for y]
4x-4x-3=-3
-3=-3 [the equations are for the same line.
.
4x-y=-3 [re-write in the slope-intercept form by solving for y]
-y=-4x-3
y=4x+3 [same as the second equation, therefore, the lines are the same]